A cone has a radius of r units and a height of h units. A cylinder has a radius of 2r units and has the same volume as the cone. WHat is the height of the cylinder?
1
Expert's answer
2011-09-06T05:49:04-0400
The volume of the cone is equal to
V = 1/3 π r2 h
The volume of the cylinder is
V = π (2r)2 H
Thus
1/3 π r2 h = π 4r2 H
1/3 h = 4 H
H = 1/12 h